OpenStudy (anonymous):

what quadrants is 4x+2y=8

OpenStudy (anonymous):

well u can use a site

OpenStudy (anonymous):

https://www.wolframalpha.com/

OpenStudy (anonymous):

it helps XD

OpenStudy (anonymous):

4x+2y=8 -4x -4x ____________________ 2y=-4x+8 /2 /2 /2 y=2x+4

OpenStudy (anonymous):

using the intercept method, let y = 0 to find out what your x-intercept is, and let x = 0 to find out what your y-intercept is. when you find the two intercepts, graph them and you will see which of the three quadrants this line will pass through

OpenStudy (anonymous):

good work gab by but use the https://www.wolframalpha.com/ to help u its mai fav cheat lol

OpenStudy (anonymous):

Ha thanks